# Understanding the RoPE Extensions of Long-Context LLMs: An Attention Perspective

### v1

@@ Abstract
LLMs, Long Context, Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) (Radford et al., 2018; Touvron et al., 2023; Zhang et al., 2023; Brown et al., 2020) have accommodated a wide range of natural language processing applications, such as code completion (Rozière et al., 2023) and question answering (Kamalloo et al., 2023; Jiang et al., 2021; Su et al., 2019). However, a notable challenge limiting further customization is possibly the inability of LLMs to utilize context beyond the pretrained length (Minaee et al., 2024; Chen et al., 2023b) due to the inherent flaw of rotary position embedding (RoPE) being used. Fortunately, RoPE extensions emerge as key ingredients to enabling LLMs to leverage extended context that exceeds pretrained scope (Chen et al., 2023b; Peng et al., 2023; Liu et al., 2023; Han et al., 2023; Rozière et al., 2023). These RoPE extensions focus on improving performance on long texts, yet frustratingly, only a few of them (Liu et al., 2023; Han et al., 2023; Men et al., 2024) have delved into the underlying mechanisms in a complicated way.
This demands us to systematically analyze common RoPE extensions more straightforwardly, from the perspective of attention (Vaswani et al., 2017). We include three widely-used RoPE extensions, i.e., position interpolation (Chen et al., 2023b), YaRN (Peng et al., 2023), and NTK-Aware interpolation (Rozière et al., 2023). To our best knowledge, there is simply no research in understanding RoPE extensions for long-context models thoroughly from an attention perspective.
As a start, we strive to primarily study these methods on a long-context perplexity test (PPL), and empirically compare their corresponding attention patterns. We discover that finetuning LLMs with these RoPE-extension methods accordingly in the pretrained length principally lifts their extrapolation performance. Particularly with the NTK-Aware interpolation method, one can extrapolate up to 32$\times$ beyond the pretrained length. To unleash the reasons behind the successes of these methods, we collect the attention scores respectively distributed in 2K and 8K lengths during inference. The results demonstrate that these methods maintain attention patterns consistent with those observed at the pretrained length. In contrast, the attention patterns of the RoPE are substantially deviated.
Afterward, following literature (Hu et al., 2024; Fu et al., 2024), we examine these RoPE extensions on a more challenging long-context test called Needle-in-a-Haystack (Needle) (Kamradt, 2023). We find that the RoPE extensions could pass more tests than the RoPE does. Nonetheless, as the context length increased, the RoPE extensions could hardly locate the needles. We associate the observation with attention uncertainty, specifically in the form of attention entropy. We uncover that large uncertainty leads to retrieval errors: the positions that incur large attention entropy are exactly where the incorrect answers are borrowed from.
@@ 2.2 RoPE and Its Extensions
where $\theta_{j}=b^{-2j/d},j\in\{0,1,...,d/2-1\}$.
Position Interpolation (PI). As described in Chen et al. (2023a) and Kaiokendev (2023), PI involves proportionally downscaling the position index $m$ to $m/\alpha$ in Equation 1.
@@
Following the default settings of the original papers (Chen et al., 2023b; Peng et al., 2023; Liu et al., 2023), we adjust $\alpha$ from 1 to 16 in $m/\alpha$ for PI and YaRN, while adjusting $b$ from 10,000 to 1,000,000 for NTK in our experiments.
@@ 2.3 Long-Context Evaluations
Following existing works (Chen et al., 2023b; Peng et al., 2023; Fu et al., 2024), we use the perplexity test (dubbed PPL) as the primary evaluation and the Needle-in-a-Haystack test as a more challenging evaluation. The perplexity is a primary measure that reflects a model’s ability to handle long texts. The Needle-in-a-Haystack test (dubbed Needle) (Kamradt, 2023) requires LLMs to accurately recall a specific sentence (the Needle) embedded at an arbitrary location within a long document (the haystack). We obtain the perplexity using the Proof-pile (Azerbayev et al., 2022) dataset. We follow the standard described in Fu et al. (2024) for the Needle-in-a-Haystack accuracy.
@@ 3 RoPE Extensions on PPL
We preliminarily study the RoPE extensions from a primary long-context perplexity test. From the test, as illustrated in Figure 1, we identify that NTK can extrapolate from 4K to 128K, whereas PI and YaRN can extrapolate to 62K. We observe similar results in both the smaller model MiniMA-2-3B and the larger model LLaMa-2-13B, as illustrated in Figures 4 and 7. To recognize why these RoPE extensions enable train-short-and-test-long properties in PPL, we collect the attention scores on 10 sequences in 2K and 8K and visualize their attention distributions. The followings are a few key takeaways from the attention perspective:
Figure 1: Perplexity on Proof-pile dataset (Lower is better).
RoPE extensions maintain the original attention patterns.
As shown in Figure 2, similar to the findings from Chen et al. (2023b), we acknowledge that the attention patterns fluctuate when the RoPE is tested on 8K sequences (exceeding the training length). However, with RoPE extensions, the attention distributions, as illustrated in Figures 2(c-e), revert to the original pattern seen in Figure 2(a) when tested on 8K sequences. Similar observations are seen in both LLaMa-2-13B and MiniMA-2-3B, as illustrated in Figures 5 and 8.
@@
Figure 2: Attention distributions of RoPE, PI, YaRN, and NTK methods on 2K and 8K sequences. The red line represents the mean attention scores across all heads, layers, and examples. The other lines indicate the attention scores for each head in each layer of the examples.
RoPE extensions closely resemble the attention patterns of models trained on longer cotnext.
To further verify whether RoPE extensions maintain the original attention patterns, we aim to directly quantify the Jensen–Shannon (JS) divergence between different attention distributions. We choose LLaMa-2 and LLaMa-3 as bases. As illustrated in the bottom row of Table 1, the JS divergence between the RoPE extensions and LLaMa-3 is smaller than between the RoPE and LLaMa-3. This indicates that the attention patterns of RoPE extensions resemble those of models directly trained in a longer context.
NTK and YaRN do not affect the attention patterns within the pretrained length.
Some RoPE extensions can degrade performance within the original pretrained length (Peng et al., 2023; Zhang et al., 2024). To verify whether RoPE extensions alter the attention patterns within the pretrained length, we also calculate the JS divergence among these models’ attention distributions at a 2K length. As illustrated on the top row of Table 1, the JS divergence for the NTK and YaRN is very low, almost zero, indicating minimal impact on attention distribution. On the contrary, the JS divergence for the PI is significantly higher. Therefore, we conclude that the NTK and YaRN methods do not affect attention patterns within the pretrained length.
(a) RoPE
(b) Finetuning with PI
(c) Finetuning with YaRN
(d) Finetuning with NTK
(e) Finetuning on 4K with NTK from (d)
(f) Finetuning on 32K with NTK from (d)
Figure 3: Performance comparison for the Needle-in-a-Haystack Test. The x-axis represents the length of the document, while the y-axis indicates the position where the needle is located. A red cell indicates that the model fails to recall the information in the needle, whereas a green cell indicates success. A white dashed line denotes the model’s continual pretrain length. Each value in the cells signifies the mean attention entropy, with higher values reflecting more dispersed attention.
4 RoPE Extensions on Needle
To understand the performance and behavior of the RoPE extensions on more challenging long-context tasks, we conduct Needle testing (Fu et al., 2024). As shown in Figure 3(a-d), LLaMa-2-7B with RoPE extensions can pass more needle tests than the RoPE. However, as the context length increases, some tests fail, resulting in retrieval needle errors. Eventually, almost all fail in extremely long contexts. We also conduct Needle testing on the MiniMA-2-3B and LLaMa-2-13B models with RoPE and PI. Unlike the LLaMa-2-7B, the PI method shows a more significant improvement in the LLaMa-2-13B, as depicted in Figure 9. In contrast, on the MiniMA-2-3B, PI passes only a few needle tests at longer lengths, as illustrated in Figure 6. We attribute these observations to the impact of model size. Below are key takeaways from the attention perspective:
Attention uncertainty leads to more retrieval needle errors.
To acquire the reason behind the retrieval needle errors, we calculate the entropy of attention for each length and depth, as illustrated in Figures 3. Our findings demonstrate that the locations of retrieval needle errors often coincide with high attention entropy. For example, at the same depth, the positions with errors are among the top-k in entropy; similarly, at the same length, the error positions also have high entropy. We hypothesize that the increase in attention entropy with longer test lengths is due to the train-short-and-test-long setting. During inference, the number of tokens handled by the self-attention mechanism far exceeds that during training. More tokens lead to more dispersed attention, i.e., higher uncertainty, causing a mismatch between training and inference.
A natural approach to lower attention uncertainty for enhancing extrapolation.
A direct solution is to train on longer contexts, thereby increasing the number of attention tokens during training and reducing attention uncertainty. To validate our hypothesis, we finetune models on 4K and 32K training lengths with the same tokens on NTK. As shown in Figures 3(e) and 3(f), compared to models trained in short contexts, models trained in more extended contexts exhibited significantly lower attention uncertainty. For example, at length 63938, the attention entropy is generally below 5. The Needle test pass rates improved significantly, especially in longer testing contexts. Conversely, models trained with the same number of tokens but shorter context sizes showed little to no change in attention entropy, remaining similar to the original one (3(d)).
5 Conclusions
Appendix A Experimental Results on MiniMA-2-3B
Figure 4: Validation Perplexity on Proof-pile (Azerbayev et al., 2022) of MiniMA-2-3B
@@ Appendix B Experimental Results on LLaMa-2-13B
Figure 7: Validation Perplexity on Proof-pile of LLaMa-2-13B

### v3

@@ Abstract
Understanding the RoPE Extensions of Long-Context LLMs:
An Attention Perspective
Meizhi Zhong1thanks: Work during Xiaohongshu internship., Chen Zhang2, Yikun Lei2, Xikai Liu2, Yan Gao2, Yao Hu2, Kehai Chen1thanks: Corresponding authors, Min Zhang1 1Institute of Computing and Intelligence, Harbin Institute of Technology, Shenzhen, China 2 Xiaohongshu Inc. meizhi.zhong.1999@gmail.com, chenzhang9702@outlook.com, {chenkehai,zhangmin2021}@hit.edu.cn, {zhizhu,xikai,yadun,xiahou}@xiaohongshu.com
@@ 1 Introduction
Large language models (LLMs) (Radford et al., 2018; Touvron et al., 2023; Zhang et al., 2023; Li et al., 2024; Zhang et al., 2024a, b) have accommodated a wide range of natural language processing applications, such as code completion (Rozière et al., 2023) and question answering (Kamalloo et al., 2023; Jiang et al., 2021; Su et al., 2019). However, a notable challenge limiting further customization is possibly the inability of LLMs to utilize context beyond the pretrained length (Minaee et al., 2024; Chen et al., 2023a) due to the inherent flaw of rotary position embedding (RoPE) being used. Fortunately, RoPE extensions emerge as key ingredients to enabling LLMs to leverage extended context that exceeds pretrained scope (Chen et al., 2023a; Peng et al., 2023; Liu et al., 2023; Han et al., 2023; Rozière et al., 2023). These RoPE extensions focus on improving performance on long texts, yet frustratingly, only a few of them Liu et al. (2023); Han et al. (2023); Men et al. (2024) have explored the underlying mechanisms in depth.
Thus, we systematically analyze common RoPE extensions more straightforwardly, from the perspective of attention (Vaswani et al., 2017). We include three widely-used RoPE extensions, i.e., position interpolation (Chen et al., 2023a), YaRN Peng et al. (2023), and NTK-Aware interpolation (Rozière et al., 2023). To our best knowledge, there is simply no research in understanding RoPE extensions for long-context models thoroughly from an attention perspective.
As a start, we strive to primarily study these methods on a long-context perplexity test (PPL) and empirically compare their corresponding attention patterns. We found that finetuning LLMs with these RoPE-extension methods which match the original pretraining length improves extrapolation performance. Particularly with the NTK-Aware interpolation method, one can extrapolate up to 32$\times$ beyond the pretrained length. To unleash the reasons behind the successes of these methods, we collect the attention scores respectively distributed in 2K and 8K lengths during inference. The results demonstrate that these methods maintain attention patterns consistent with those observed at the pretrained length. In contrast, the attention patterns of the RoPE are substantially deviated.
Afterward, following literature (Fu et al., 2024), we examine these RoPE extensions on a more challenging long-context test called Needle-in-a-Haystack (Needle) (Kamradt, 2023). We find that the RoPE extensions could pass more tests than the RoPE does. Nonetheless, as the context length increased, the RoPE extensions could hardly locate the needles. We associate the observation with attention uncertainty. We uncover that large uncertainty leads to retrieval errors: the positions that incur large attention uncertainty are exactly where the incorrect answers are borrowed from.
@@ 2.2 RoPE and Its Extensions
where $\theta_{j}=b^{-2j/d},j\in\{0,1,...,d/2-1\}$, and $b$ represents the base frequency for RoPE.
Position Interpolation (PI). As described in Chen et al. (2023b) and Kaiokendev (2023), PI involves proportionally downscaling the position index $m$ to $m/\alpha$ in Equation 1.
@@
Following the default settings of the original papers (Chen et al., 2023a; Peng et al., 2023; Liu et al., 2023), we adjust $\alpha$ from 1 to 16 in $m/\alpha$ for PI and YaRN, while adjusting $b$ from 10,000 to 1,000,000 for NTK in our experiments.
@@ 2.3 Long-Context Evaluations
Following existing works (Chen et al., 2023a; Peng et al., 2023; Fu et al., 2024), we use the perplexity test (dubbed PPL) as the primary evaluation and the Needle-in-a-Haystack test as a more challenging evaluation. The perplexity is a primary measure that reflects a model’s ability to handle long texts. The Needle-in-a-Haystack test (dubbed Needle) (Kamradt, 2023) requires LLMs to accurately recall a specific sentence (the Needle) embedded at an arbitrary location within a long document (the haystack). We obtain the perplexity on the Proof-pile (Azerbayev et al., 2022) dataset. We follow the standard described in Fu et al. (2024) for the Needle-in-a-Haystack accuracy.
(a) LLaMa-2-7B
(b) MiniMA-2-3B
(c) LLaMa-2-13B
Figure 1: Perplexity on Proof-pile(Lower is better).
@@ 3 RoPE Extensions on PPL
We study RoPE extensions by comparing performance on long-context perplexity testing. From the test, as illustrated in Figure 1, we identify that NTK can extrapolate from 4K to 128K, whereas PI and YaRN can extrapolate to 62K. We observe similar results in both the smaller model MiniMA-2-3B and the larger model LLaMa-2-13B, as illustrated in Figures 1(b) and 1(c). To recognize why these RoPE extensions enable train-short-and-test-long properties in PPL, we collect the attention scores on 10 sequences in 2K and 8K and visualize their attention distributions. The followings are a few key takeaways from the attention perspective:
RoPE extensions maintain the original attention patterns. As shown in Figure 2, similar to the findings from Chen et al. (2023a), we observe that the attention patterns fluctuate when the RoPE is tested on 8K sequences (exceeding the training length). However, with RoPE extensions, the attention distributions, as illustrated in Figures 2(c-e), revert to the original pattern seen in Figure 2(a) when tested on 8K sequences. Similar observations are seen in both LLaMa-2-13B and MiniMA-2-3B, as illustrated in Figures 3 and 6.
@@
Figure 2: Attention distributions of RoPE, PI, YaRN, and NTK methods on 2K and 8K sequences. The red line represents the mean attention scores across all heads, layers, and examples. The other lines indicate the attention scores for each head in each layer.
@@
(c) Finetuning with YaRN
(d) Finetuning(FT) with NTK
(e) FT on 4K with NTK from (d)
(f) FT on 32K with NTK from (d)
Figure 4: Performance comparison for the Needle-in-a-Haystack Test. The x-axis represents the length of the document, while y-axis indicates the depth percentage, showing the needle’s position within the document. For instance, a position of 50% signifies that the needle is placed in the middle of the document. A red cell indicates that the model fails to recall the information in the needle, whereas a green cell indicates success. A white dashed line denotes the model’s continual pretrain length. Each value in the cells signifies the mean attention entropy, with higher values reflecting more dispersed attention.
RoPE extensions closely resemble the attention patterns of models trained on longer context. To further verify whether RoPE extensions maintain the original attention patterns, we aim to directly quantify the Jensen–Shannon (JS) divergence between different attention distributions. Using LLaMa-2 and LLaMa-3 as baselines, we collected 10,240 samples of attention distributions to calculate the JS divergence. As illustrated in the bottom row of Table 1, the JS divergence between the RoPE extensions and LLaMa-3 is more minor than between the RoPE and LLaMa-3. This indicates that the attention patterns of RoPE extensions resemble those of models directly trained on a longer context.
NTK and YaRN do not affect the attention patterns within the pretrained length. Some RoPE extensions can degrade performance within the original pretrained length (Peng et al., 2023; Zhang et al., 2024c). To verify whether RoPE extensions alter the attention patterns within the pretrained length, we also calculate the JS divergence among these models’ attention distributions at a 2K length. As illustrated on the top row of Table 1, the JS divergence for the NTK and YaRN is very low, almost zero, indicating minimal impact on attention distribution. On the contrary, the JS divergence for the PI is significantly higher. Therefore, we conclude that the NTK and YaRN methods do not affect attention patterns within the pretrained length.
4 RoPE Extensions on Needle
To understand the performance and behavior of the RoPE extensions on more challenging long-context tasks, we conduct Needle testing Fu et al. (2024). As shown in Figure 4(a-d), LLaMa-2-7B with RoPE extensions can pass more needle tests than the RoPE. However, as the context length increases, some tests fail, resulting in needle retrieval errors. Eventually, almost all fail in extremely long contexts. We also conduct Needle testing on the MiniMA-2-3B and LLaMa-2-13B models with RoPE and PI. Unlike the LLaMa-2-7B, the PI method shows a more significant improvement in the LLaMa-2-13B, as depicted in Figure 7. In contrast, on the MiniMA-2-3B, PI passes only a few needle tests at longer lengths, as illustrated in Figure 5. We attribute these observations to the impact of model size. Below are key takeaways from the attention perspective:
Attention uncertainty leads to more needle retrieval errors. To find the reason behind the needle retrieval errors, we calculate the entropy of attention for each length and depth, as illustrated in Figures 4. For details on the calculation of attention entropy, please refer to Appendix C. Our findings demonstrate that the locations of needle retrieval errors often coincide with high attention entropy. For example, at the same depth, the positions with errors are among the top-k in entropy; similarly, at the same length, the error positions also have high entropy. We hypothesize that the increase in attention entropy with longer test lengths is due to the train-short-and-test-long setting. During inference, the number of tokens handled by the self-attention mechanism far exceeds that during training. More tokens lead to more dispersed attention, i.e., higher uncertainty, causing a mismatch between training and inference.
A natural approach to lower attention uncertainty for enhancing extrapolation. A direct solution is to train on longer contexts, thereby increasing the number of attention tokens during training and reducing attention uncertainty. To validate our hypothesis, we finetune models on 4K and 32K training lengths with the same tokens on NTK. As shown in Figures 4(e) and 4(f), compared to models trained in short contexts, models trained in more extended contexts exhibited significantly lower attention uncertainty. For example, at length 63938, the attention entropy is generally below 5. The Needle test pass rates improved significantly, especially in longer testing contexts. Conversely, models trained with the same number of tokens but shorter context sizes showed little to no change in attention entropy, remaining similar to the original one (4(d)).
5 Conclusions
Limitations
This paper primarily analyzes the widely-used decoder-only LM, LLaMa (Touvron et al., 2023). It does not include a validation study of encoder-decoder and encoder-only architectures.
Acknowledgements
We would like to thank the anonymous reviewers and meta-reviewer for their insightful suggestions. This work was supported by the National Natural Science Foundation of China under Grant U23B2055 and 62276077, and Shenzhen Science and Technology Program under Grant ZDSYS20230626091203008.
Appendix A Experimental Results on MiniMA-2-3B
(a) RoPE
(b) Finetuning with PI
@@ Appendix A Experimental Results on MiniMA-2-3B
Similar to the analysis in § 4, the Needle-in-a-Haystack Test for MiniMa-2-3B also indicates that the locations of needle retrieval errors frequently align with areas of high attention entropy.
@@ Appendix B Experimental Results on LLaMa-2-13B
Consistent with the analysis in § 3, we observe that the attention patterns fluctuate when RoPE is applied to 8K sequences, which exceed the training length. However, when using RoPE extensions, the attention distributions return to their original patterns for 8K sequences, as demonstrated in Figures 6.
@@
Similar to the analysis in § 4, the Needle-in-a-Haystack Test for LLaMa-2-13B also indicates that the locations of needle retrieval errors frequently align with areas of high attention entropy.
Appendix C Detailed Calculation of Attention Entropy
Algorithm 1 Calculation of Attention Entropy
1:Input: model, prompt
2:Output: average attention entropy score
3:procedure AttentionEntropy($model,prompt$)
4:     Initialize $entropy\_list\leftarrow[]$
5:     $output\_tokens\leftarrow\text{[ ]}$
6:     while not end of generation do
7:         $token,attention\_distribution\leftarrow\text{GenerateTokenAndGetAttention}(model,prompt\,+\,output\_tokens)$
8:         $output\_tokens\text{.append}(token)$
9:         $entropy\leftarrow\text{CalculateEntropy}(attention\_distribution)$
10:         $entropy\_list\text{.append}(entropy)$
11:     end while
12:     $average\_entropy\leftarrow\text{Average}(entropy\_list)$
13:     return $average\_entropy$
14:end procedure
15:function CalculateEntropy($distribution$)
16:     $entropy\leftarrow 0$
17:     for all $p\text{ in }distribution$ do
18:         if $p>0$ then
19:              $entropy\leftarrow entropy-p\log(p)$
20:         end if
21:     end for
22:     return $entropy$
23:end function